(1-Dimensional Collision) A 100g solid ball approaches a 1kg solid ball which is also approaching the 100g ball. If the initial speeds of both balls are 5m/s, what are the final speeds of the balls?


1
Expert's answer
2021-04-26T09:48:36-0400

To be given in question

Motion 1D

M1=100gm =0.1kg

M2=1000gm=1.0kg

u1=u2=5 meter/sec

To be asked in question

v1=?

v2=?

We know that

Momentum conservation law

"m_{1}u_{1}+m_{2}u_{2} =m_{1}v_{1}+m_{2}v_{2} \rightarrow(1)"

Put value in eqution (1)

"0.1\times5+1\times5=0.1\times v_{1}+1\times v_{2}"

0.5+5.0="(0.1v_{1}+1v_{2})"

5.5="(0.1v_{1}+1v_{2})\rightarrow(2)"

We know that

Cofficient of restitution

"e=1"

elastic collision

"v_{1}-v_{2} =u_{2}-u_{1}\rightarrow(3)"

Put value in eqution (3)

"v_{1}-v_{2}=0"

"v_{1}=v_{2}"

equation (2)put values "v_{1}andv_{2}"

"5.5=1.1v_{1}"

"v_{1}=\frac{55}{11}"

"v_{1}=5meter \/sec"

Similarly

"v_{1}=v_{2}=5meter\/sec"